If the hydroxide ion concentration in solution is 2.648 x 10-8 M, what is the pH of the solution? Assume the solution is at 25 ° C. Provide your response to one digit after the decimal.
Answer #1

pOH = -log[OH-]
pOH = -log(2.648*10-8)
= 7.577
pH = 14-pOH
= 14-7.577
= 6.423
= 6.4

Answer : 6.4